An airplane lands going 80 m/s. It stops in 16 seconds. What is its acceleration? How far does it go while stopping?
Answer #1

_____________________________________________________

Initial velocity u = 80 m / s
final velocity v = 0
time t= 16 s
accleration a = ( v - u ) / t = ( 0 - 80 ) / 16 = - 5 m/s2
_____________________________________________________
from the relation v ^ 2- u^ 2= 2aS
required distance S = [ v^ 2- u^ 2] / 2a = ( 0 - 802 )/ 2* (-5) = 640 m
_____________________________________________________

Please rate
if any mistake in this answer please comment i will clarify your doubt . thank you